# Improving LLM Unlearning Robustness via Random Perturbations

### v3

@@ Appendix D Robustness of RNA Against Hard Negative Forget-Tokens
In this section, we analyze RNA’s robustness to forget-tokens that exhibit high $n$-gram similarity.
@@
Setup.
For each document in the WMDP forget-set, we extract $n$-grams for $n\in\{2,4,8,16\}$ and compute their feature embeddings using Sentence-BERT [47], along with the embedding of the full document. We then extract the top $10$ most similar $n$-grams to each document based on embedding cosine similarity. Perturbed MMLU QAs corresponding to these $n$-gram forget-tokens are synthesized following the procedure outlined in Subsection A.2.
Hyperparameters.
We utilize model checkpoints from the previous setting and perform evaluations accordingly. Results are reported for checkpoints selected based on the optimal noise scale $\nu$, as detailed in Table 1.
@@
w/ RNA $32.2$ ($-5.1$) $56.7$ ($+0.9$) $44.4$ ($+11.2$)
@@ Results.
RNA’s performance is summarized in Table 3 through Table 5. We observe that RNA consistently improves the robustness of unlearned models across all $n$-gram perturbations. The most pronounced gains are observed when RNA is applied with MSE retain-losses. Specifically, for DPO+MSE, performance improvements are $+24.6$ ($2$-gram), $+29.5$ ($4$-gram), $+24.9$ ($8$-gram), and $+19.6$ ($16$-gram); for SimNPO+MSE, gains are $+25.4$, $+23.3$, $+16.5$, and $+11.2$. Importantly, RNA introduces minimal impacts on MMLU performance, where changes are generally within less than $0.5$. However, RNA tends to reduce WMDP accuracy across all methods, with slightly drop ranging from $0.5$ to $5.0$. Additionally, RM methods derive minimal benefit from RNA under these settings.
@@
Figures 810 demonstrate that RNA generally improves the robustness of unlearned models. While Figures 8 and 9 show improvements in both settings, no consistent trend emerges across all methods. Notably, models trained with MSE retain-loss achieve significant gains from RNA. Figure 10 further shows that injecting noise into all layer is particularly effective at moderate noise levels (e.g., $1\times 10^{-3}$). However, as the noise scale $\nu$ increases, model accuracy declines sharply. Importantly, MMLU accuracy remains stable with RNA integration, highlighting that RNA not only boosts robustness but also preserves general knowledge and capabilities.
Appendix F Limitation and Future Work
@@
Limitations.
Due to computational constraints, experiments are conducted only on the 7B model and with updates to a limited set of layer parameters, which may risk overlooking interesting aspects of RNA’s generalization. Although RNA has demonstrated effectiveness, it relies heavily on hyperparameter grid search to identify an optimal noise scale makes it impractical for extremely large models with hundreds of billions of parameters.
Future works.
RNA is designed for fine-tuning-based unlearning, which requires access to the model parameters. Future work exploring RNA as an inference-time intervention method could be promising. One potential direction is to conduct analysis on the optimal coefficient or noise scale to further improve RNA’s effectiveness.
Appendix G Broader Impact
We establish a novel theoretical framework that bridges the connection between machine unlearning and backdoor attacks, providing crucial insights into the vulnerabilities of unlearned models. Our theoretical and empirical analysis providing a valuable solution for developing more secure and reliable machine unlearning systems.

### v4

@@ Appendix D Robustness of RNA Against Hard Negative Forget-Tokens
Analysis on the harmfulness of forget-tokens. One might ask: “Which forget-tokens when appearing in the retain-query can cause the unlearned model to misbehave?”. We examine the harmfulness of forget-tokens in the forget-set by measuring the cosine similarity between bi-gram forget-tokens and their respective documents, across all documents in the WMDP forget-sets. We select the top $10$ most similar, least similar, and those with values around the mean of the distribution. Perturbed MMLU QAs with respect to these forget-tokens are synthesized following the procedure described in Section 6.1. As shown in Figure 10, we observed a clear trend between the accuracy and the similarity: forget-tokens with higher similarity with their corresponding documents are more harmful to unlearned models. We further assess the RNA models’ robustness against $n$-gram similarity perturbations for $n\in\{4,8,16\}$.
Figure 10: Accuracy of unlearned models on perturbed MMLU with respect to bi-gram similarity perturbations.
@@
w/ RNA $32.2$ (5.1-5.1) $56.7$ (+0.9+0.9) $44.4$ (+11.2+11.2)
Setup.
For each document in the WMDP forget-set, we extract $n$-grams for $n\in\{2,4,8,16\}$ and compute their feature embeddings using Sentence-BERT (Reimers & Gurevych, 2019), along with the embedding of the full document. We then extract the top $10$ most similar $n$-grams to each document based on embedding cosine similarity. Perturbed MMLU QAs corresponding to these $n$-gram forget-tokens are synthesized following the procedure outlined in Subsection A.2. We utilize model checkpoints from the previous setting and perform evaluations accordingly. Results are reported for checkpoints selected based on the optimal noise scale $\nu$, as detailed in Table 1.
@@ Results.
RNA’s performance is summarized in Table 3 through Table 5. We observe that RNA consistently improves the robustness of unlearned models across all $n$-gram perturbations. The most pronounced gains are observed when RNA is applied with MSE retain-losses. Specifically, for DPO+MSE, performance improvements are +24.6+24.6 (22-gram), +29.5+29.5 (44-gram), +24.9+24.9 (88-gram), and +19.6+19.6 ($16$-gram); for SimNPO+MSE, gains are +25.4+25.4, +23.3+23.3, +16.5+16.5, and +11.2+11.2. Importantly, RNA introduces minimal impacts on MMLU performance, where changes are generally within less than $0.5$. However, RNA tends to reduce WMDP accuracy across all methods, with slightly drop ranging from $0.5$ to $5.0$. Additionally, RM methods derive minimal benefit from RNA under these settings.
@@
Figure 1113 demonstrate that RNA generally improves the robustness of unlearned models. While Figure 11 and 12 show improvements in both settings, no consistent trend emerges across all methods. Notably, models trained with MSE retain-loss achieve significant gains from RNA. Figure 13 further shows that injecting noise into all layers is particularly effective at moderate noise levels (e.g., $1\times 10^{-3}$). However, as the noise scale $\nu$ increases, model accuracy declines sharply. Importantly, MMLU accuracy remains stable with RNA integration, highlighting that RNA not only boosts robustness but also preserves general knowledge and capabilities.
Appendix F Robustness of Unlearned Models Against Prompt Attacks
The retaining process is reframed as a backdoor defense against a specific type of backdoor (forget-tokens). The noise injection in RNA is reminiscent of adversarial training. There is a well-known phenomenon that, when defending against one type of attack, can inadvertently create new vulnerabilities or increase susceptibility to other attacks (Tramer & Boneh, 2019; Weng et al., 2020; Kamath et al., 2021) on the general capabilities. In this section, we present an analysis of whether RNA makes the model become more susceptible to other adversarial attacks.
Setup.
We employ four widely used adversarial attack methods to evaluate the side effects of RNA, including Greedy Coordinate Gradient (GCG; Zou et al. (2023)), TextBugger (Li et al., 2018), DeepWordBug (Gao et al., 2018), and TextFooler (Jin et al., 2020). TextFooler is an adversarial word-substitution method that relies on importance scores to identify and replace important words with corresponding synonyms. TextBugger generates adversarial prompts by augmenting text through character-level perturbations. DeepWordBug generates adversarial prompts by introducing character-level perturbations, such as insertions, deletions, and substitutions. GCG is a gradient-based attack that iteratively modifies injected adversarial tokens along the directions with respect to the largest increase of loss to construct adversarial prompts. We utilize optimal checkpoints from the main setting, as detailed in Subsection A.4.
Empirical effects of unlearning against attacks.
We report the accuracy and ROUGE-L score under attack of the original unlearned model and RNA models in Table 6. As we can observe in this table, under attacks, all unlearning methods consistently reduce models’ robustness, making the models more vulnerable to adversarial prompt attacks. For instance, the base model achieves $40.3$ AuA under GCG attack, whereas unlearned models drop to the range of $30\to 39$ (e.g., RMU $33.6$, Adaptive RMU $38.5$, RSV $39.2$, NPO+KL $35.4$). Similar reductions are observed under TextBugger (base $33.6$ vs. unlearned $26\to 31$) and DeepWordBug (base $39.6$ vs. unlearned $28\to 38$, except NPO+MSE: $40.3$).
Effects of RNA on model robustness.
We observed that RNA’s impact is dependent on the underlying unlearning method, with no clear trends observed. In summary, unlearning uniformly reduces models’ robustness, while RNA can partially mitigate these vulnerabilities in some cases.
Methods GCG TextBugger DeepWordBug TextFooler
AuA ROUGE-L AuA ROUGE-L AuA ROUGE-L AuA ROUGE-L
Base Original $40.3$ $33.6$ $39.6$ $52.9$
Representation Misdirection
RMU Original $33.6$ $63.0$ $30.5$ $81.2$ $38.2$ $76.8$ $50.5$ $85.2$
w/ RNA 40.340.3+6.7+6.7 60.960.92.1-2.1 30.530.5+0.0+0.0 79.179.11.1-1.1 38.938.9+0.7+0.7 76.476.40.4-0.4 50.150.10.4-0.4 84.284.21.0-1.0
Adap. RMU Original $38.5$ $63.5$ $30.5$ $80.0$ $38.9$ $76.2$ $49.8$ $83.3$
w/ RNA 43.543.5+5.0+5.0 62.462.41.1-1.1 30.830.8+0.3+0.3 75.275.24.8-4.8 39.239.2+0.3+0.3 72.772.73.5-3.5 50.850.8+1.0+1.0 79.279.24.1-4.1
RSV Original $39.2$ $63.0$ $30.8$ $78.2$ $38.5$ $75.3$ $48.7$ $80.6$
w/ RNA 38.238.21.0-1.0 62.562.50.5-0.5 27.027.03.8-3.8 77.177.11.1-1.1 35.035.03.5-3.5 73.773.71.6-1.6 44.244.24.5-4.5 83.283.2+2.6+2.6
Preference Optimization
NPO+KL Original $35.4$ $51.2$ $26.3$ $61.3$ $34.7$ $61.3$ $43.1$ $67.7$
w/ RNA 31.231.24.2-4.2 55.055.0+4.8+4.8 25.225.21.1-1.1 64.164.1+2.8+2.8 31.231.23.5-3.5 62.362.3+1.0+1.0 39.239.23.9-3.9 67.067.00.7-0.7
NPO+MSE Original $40.7$ $57.6$ $29.1$ $66.7$ $40.3$ $66.0$ $46.6$ $71.2$
w/ RNA 34.334.36.4-6.4 51.651.66.0-6.0 24.224.24.9-4.9 66.266.20.5-0.5 34.334.36.0-6.0 64.464.41.6-1.6 46.346.30.3-0.3 70.970.90.3-0.3
DPO+KL Original $30.1$ $47.3$ $27.7$ $58.1$ $34.0$ $57.3$ $41.7$ $61.1$
w/ RNA 29.829.80.3-0.3 56.456.4+9.1+9.1 29.129.1+1.4+1.4 67.067.0+8.9+8.9 33.333.30.7-0.7 64.764.7+7.4+7.4 42.842.8+1.1+1.1 68.968.9+7.8+7.8
DPO+MSE Original $28.0$ $50.2$ $19.2$ $61.1$ $28.4$ $61.5$ $36.8$ $65.9$
w/ RNA 27.727.70.3-0.3 55.655.6+5.4+5.4 23.523.5+4.3+4.3 57.557.53.6-3.6 30.530.5+2.1+2.1 58.458.43.1-3.1 39.639.6+2.8+2.8 64.364.31.6-1.6
SimNPO+KL Original $29.1$ $49.9$ $27.0$ $61.1$ $34.7$ $60.7$ $41.7$ $64.2$
w/ RNA 30.130.1+1.0+1.0 55.055.0+5.1+5.1 27.027.0+0.0+0.0 61.761.7+0.6+0.6 35.435.4+0.7+0.7 60.860.8+0.1+0.1 44.244.2+2.5+2.5 67.067.0+2.8+2.8
SimNPO+MSE Original $35.4$ $52.3$ $29.8$ $66.7$ $36.8$ $66.8$ $44.5$ $72.3$
w/ RNA 38.238.2+3.8+3.8 58.658.6+6.3+6.3 31.531.5+1.7+1.7 71.971.9+5.2+5.2 42.842.8+6.0+6.0 70.170.1+3.3+3.3 48.448.4+3.9+3.9 75.775.7+3.4+3.4
Table 6: Accuracy under attack (AuA $\uparrow$) and ROUGE-L ($\uparrow$) of unlearning methods on adversarial perturbed MMLU, comparing Original vs. w/ RNA. Improvements are shown in blue, drops in red.
Appendix G Effects of RNA on Model Alignment
Methods TruthfulQA ToxiGen WinoGrande CommonsenseQA HellaSwag ARC E. ARC C. BoolQ
Base Original $38.5$ $45.2$ $72.3$ $66.1$ $63.9$ $81.2$ $57.0$ $84.9$
Representation Misdirection
RMU Original $38.6$ $45.1$ $72.8$ $65.3$ $63.7$ $80.6$ $56.3$ $84.5$
w/ RNA 37.837.80.8-0.8 44.344.30.8-0.8 72.472.40.4-0.4 65.565.5+0.2+0.2 63.763.7+0.0+0.0 80.380.30.3-0.3 55.955.90.4-0.4 84.384.30.2-0.2
Adap. RMU Original $38.6$ $45.5$ $72.3$ $65.7$ $63.6$ $80.8$ $55.6$ $84.5$
w/ RNA 38.838.8+0.2+0.2 44.144.11.4-1.4 73.073.0+0.7+0.7 65.865.8+0.1+0.1 63.563.50.1-0.1 80.380.30.5-0.5 56.156.1+0.5+0.5 84.484.40.1-0.1
RSV Original $39.6$ $46.0$ $72.3$ $64.4$ $63.6$ $80.6$ $56.8$ $84.5$
w/ RNA 39.539.50.1-0.1 45.445.40.6-0.6 71.671.60.7-0.7 64.564.5+0.1+0.1 63.363.30.3-0.3 80.380.30.3-0.3 56.456.40.4-0.4 84.384.30.2-0.2
Preference Optimization
NPO+KL Original $42.9$ $45.0$ $70.8$ $64.1$ $61.8$ $80.0$ $56.7$ $84.7$
w/ RNA 40.340.32.6-2.6 44.544.50.5-0.5 70.770.70.1-0.1 62.962.91.2-1.2 61.861.8+0.0+0.0 80.280.2+0.2+0.2 56.556.50.2-0.2 84.684.60.1-0.1
NPO+MSE Original $37.3$ $45.1$ $72.1$ $62.7$ $63.0$ $80.6$ $56.4$ $85.2$
w/ RNA 36.336.31.0-1.0 44.744.70.4-0.4 72.372.3+0.2+0.2 62.262.20.5-0.5 62.962.90.1-0.1 80.780.7+0.1+0.1 56.456.4+0.0+0.0 85.285.2+0.0+0.0
DPO+KL Original $39.5$ $45.9$ $71.8$ $62.0$ $61.4$ $79.4$ $55.1$ $84.4$
w/ RNA 38.038.01.5-1.5 43.943.92.0-2.0 69.769.72.1-2.1 62.362.3+0.3+0.3 61.661.6+0.2+0.2 79.679.6+0.2+0.2 55.755.7+0.6+0.6 84.884.8+0.4+0.4
DPO+MSE Original $34.2$ $44.7$ $72.1$ $56.7$ $62.3$ $78.9$ $54.1$ $84.3$
w/ RNA 32.132.12.1-2.1 42.542.52.2-2.2 71.671.60.5-0.5 57.257.2+0.5+0.5 61.661.60.7-0.7 78.878.80.1-0.1 53.453.40.7-0.7 84.484.4+0.1+0.1
SimNPO+KL Original $43.3$ $44.0$ $71.5$ $63.8$ $62.0$ $80.0$ $56.4$ $84.5$
w/ RNA 42.142.11.2-1.2 45.245.2+1.2+1.2 71.171.10.4-0.4 61.961.91.9-1.9 61.961.90.1-0.1 79.979.90.1-0.1 56.456.4+0.0+0.0 84.884.8+0.3+0.3
SimNPO+MSE Original $38.3$ $45.8$ $71.5$ $63.8$ $62.9$ $80.6$ $56.8$ $85.6$
w/ RNA 38.338.3+0.0+0.0 43.443.42.4-2.4 70.670.60.9-0.9 62.362.31.5-1.5 62.562.50.4-0.4 80.380.30.3-0.3 56.356.30.5-0.5 85.285.20.4-0.4
Table 7: Performance of unlearning methods on 88 tasks, comparing Original vs. w/ RNA. Improvements are shown in blue, drops in red.
PO unlearning methods like DPO are themselves alignment techniques. RNA enhances robustness by increasing the diversity (via random noise) of the retain-representations. This could create potential conflicts with the precision of model alignment. We evaluate the RNA’s effect on the model’s alignment, such as faithfulness and hallucination on TruthfulQA (Lin et al., 2022) (multiple-choice QA) and ToxiGen (Hartvigsen et al., 2022), commonsense reasoning on WinoGrande (Sakaguchi et al., 2021) and CommonsenseQA (Talmor et al., 2019), natural language inference/completion on HellaSwag (Zellers et al., 2019), science reasoning on ARC (Clark et al., 2018) (easy and challenge), and factuality on BoolQ (Clark et al., 2019). Performance of unlearning methods on these tasks is shown in Table 7. As we can observe that RNA does not systematically degrade alignment, and overall alignment performance remains stable.
Appendix H Performance of Other Models
Our experiments in the main text are based on the Zephyr-7B model, which serves as a representative setup. To assess RNA’s generalization beyond the original setup, we conducted additional experiments using the Llama-3-8B (Dubey et al., 2024) and Mistral-7B Jiang et al. (2023) models on two representative unlearning methods from 2 classes: RMU and NPO+KL, across various tasks. These results provide further empirical evidence for RNA’s generalization and robustness.
Hyperparameters.
All models are fine-tuned using AdamW (Loshchilov & Hutter, 2019) for $T=500$ update steps with a learning rate of $5\times 10^{-5}$, batch size $4$, and maximum sequence length of $500$ for WMDP-Biology and $768$ for WMDP-Cyber. The unlearned layer is fixed at $l=7$. Retain weights are set to $\alpha_{\text{biology}}=\alpha_{\text{cyber}}=1200$ for both models. The coefficient values are $c_{\text{biology}}=c_{\text{cyber}}=20$ for Llama-3-8B and $c_{\text{biology}}=c_{\text{cyber}}=6.5$ for Mistral-7B. For NPO+KL, we perform a grid search over $(\alpha_{\text{biology}},\alpha_{\text{cyber}})$ and select $(5,10)$ for Llama-3-8B and $(30,40)$ for Mistral-7B. For RMU w/ RNA, we set the perturbed layer to $l=7$, tune the noise scale via grid search, and report the best performance at $\nu=7\times 10^{-2}$ (Llama-3-8B) and $\nu=3\times 10^{-2}$ (Mistral-7B). For NPO+KL w/ RNA, we also perturb layer $l=7$ and select the best scales $\nu=6\times 10^{-2}$ (Llama-3-8B) and $\nu=3\times 10^{-3}$ (Mistral-7B).
Results.
As shown in Table 8, across Llama-3-8B and Mistral-7B, RNA significantly enhances unlearning robustness of models while introducing a small trade-off on forget performance. For forget-tasks (WMDP-Biology and Cyber), RNA slightly increases the accuracy, e.g., RMU on Llama-3-8B drops from $31.4\to 34.6$ (3.2-3.2) and NPO+KL from $27.9\to 33.2$ (5.3-5.3). For retain-tasks (MMLU and Perturbed MMLU and MMLU subsets), RNA substantially improves performance, particularly on perturbed MMLU and MMLU subsets (C. Bio. and C. Sec.). For example, perturbed MMLU on Llama-3-8B with NPO+KL improves from $26.2\to 47.3$ (+21.1+21.1), and MMLU C. Bio. from $30.5\to 55.5$ (+25.0+25.0). Overall, RNA effectively recovers or enhances accuracy on retain-tasks while slightly compromising forget-task performance, demonstrating a favorable trade-off between unlearning and model robustness.
Models & Methods Llama-3-8B Mistral-7B
RMU NPO+KL RMU NPO+KL
Orig. w/ RNA Orig. w/ RNA Orig. w/ RNA Orig. w/ RNA
WMDP ($\downarrow$) $31.4$ 34.634.63.2-3.2 $27.9$ 33.233.25.3-5.3 $31.7$ 31.731.7+0.0+0.0 $29.3$ 34.034.04.7-4.7
MMLU ($\uparrow$) $60.3$ 60.260.20.1-0.1 $53.8$ 54.454.4+0.6+0.6 $58.2$ 58.658.6+0.4+0.4 $56.5$ 56.456.40.1-0.1
Perturbed MMLU ($\uparrow$) $34.4$ 47.347.3+12.9+12.9 $26.2$ 47.347.3+21.1+21.1 $27.2$ 42.242.2+15.0+15.0 $31.4$ 53.553.5+22.1+22.1
MMLU C. Bio. ($\uparrow$) $34.7$ 60.460.4+25.7+25.7 $30.5$ 55.555.5+25.0+25.0 $25.0$ 38.138.1+13.1+13.1 $32.6$ 52.052.0+19.4+19.4
MMLU C. Sec. ($\uparrow$) $29.0$ 33.033.0+4.0+4.0 $30.0$ 46.046.0+16.0+16.0 $33.0$ 46.046.0+13.0+13.0 $35.0$ 30.030.05.0-5.0
Table 8: Performance of Llama-3-8B and Mistral-7B on WMDP and MMLU, Perturbed MMLU, MMLU subsets benchmarks using RMU and NPO+KL, comparing Original vs. w/ RNA. Improvements are shown in blue, drops in red.
Appendix I Limitation and Future Work
@@
We have evaluated our methods primarily on WMDP, a widely used and representative benchmark. We acknowledge the existence of other benchmarks, such as MUSE (Shi et al., 2025) and TOFU (Maini et al., 2024), but these are less suitable for our experimental setup. Specifically, TOFU is designed to remove the influence of specific data points, making it less applicable in generative settings. While MUSE could be suitable, previous work (Shi et al., 2025) has shown that methods evaluated on MUSE often exhibit over-forgetting or under-forgetting. Since our study focuses on retain-robustness, which requires a careful balance between forgetting and retaining, MUSE is not ideal. These factors make it challenging to directly apply MUSE and TOFU in our current experiments.
Due to computational constraints, experiments are conducted only on the 7B or 8B models and with updates to a limited set of layer parameters, which may risk overlooking interesting aspects of generalization. Although RNA has demonstrated effectiveness, it relies heavily on hyperparameter grid search to identify an optimal noise scale, making it computationally expensive for extremely large models with hundreds of billions of parameters.
RNA is designed for fine-tuning-based unlearning, which requires access to the model parameters. Future work exploring RNA as an inference-time intervention method could be promising. One potential direction is to analyze the optimal coefficient or noise scale to further improve RNA’s effectiveness.
Appendix J Broader Impact
We establish a novel theoretical framework that bridges the connection between machine unlearning and backdoor attacks, providing crucial insights into the vulnerabilities of unlearned models. Our theoretical and empirical analysis provides a valuable solution for developing more secure and reliable machine learning systems.
Appendix K AI Usage Declaration
AI tools were used for grammar checking and formatting of tables and figures. All technical content and implementations were written by the authors.